OpenStudy (anonymous):

The value of a game console decreases by 60% each year. If you purchased this game console today for $150, how much would it be worth 2 years later? Please help me work through this problem.

OpenStudy (anonymous):

after 1st year 60\[60/100 * 150 =90\] then after 2nd year \[40/100*90 =24\] 24 is the answer finally

OpenStudy (anonymous):

\[150(.4)^2\]after two years it would be worth

OpenStudy (anonymous):

reason as follows. since it decreases by 60% each year, it retains 40% of its value. 40% of 150 for the first year, then 40% of that amount for the next year. year one: \[150\times .4=60\] year two \[60\times .4=24\] or we could get this directly via \[150(.4)^2=24\]